A PARTICLE STARTS FROM REST AND MOVES UNDER CONSTANT ACCELERATION IN A STRAIGHT LINE . IF TIME TAKEN BY IT TO COVER DISPLACEMENT d is 10 sec , THEN THE TIME TAKEN BY ITR COVER NEXT d DISPLACEMENT IS WHAT ?

A PARTICLE  STARTS FROM REST AND MOVES UNDER CONSTANT ACCELERATION IN A STRAIGHT LINE . IF TIME TAKEN BY IT TO COVER DISPLACEMENT d  is 10 sec , THEN THE TIME TAKEN BY ITR COVER NEXT d DISPLACEMENT IS WHAT ?

Grade:12

1 Answers

nishant narkar
5 years ago
S=ut+1/2 at^2
d=o+1/2a100
d=50 a
Velocity of object aftr completion 10 sec
V= u+at
V=a 10
For Next d diatance
50a = 10a t + 1/2 a t^2
Please solve the equation you will get time t